# Map2Text: New Content Generation from Low-Dimensional Visualizations

### v1

@@ 5. Related Work
Low-Dimensional Visualization of Text Corpora
@@
Low-dimensional data visualization is a versatile tool used across numerous fields to provide intuitive representations of large datasets (Sadiku et al., 2016). In public health, it assists in tracking disease trends and guiding data-driven decision-making (Kilimba et al., 2015). In renewable energy, visualization helps optimize energy production and consumption (Kumar and Goyal, 2016), while environmental science relies on geographic visualization to interpret ecological data (Bohman et al., 2015; MacEachren and Taylor, 2013). Fraud detection benefits from interactive visualizations that help identify suspicious patterns (Dilla and Raschke, 2015), and library management uses visualization techniques to manage resources and support strategic decisions (Murphy, 2013). A variety of methods are employed to achieve these visualizations. Textual embedding models (Muennighoff et al., 2022), such as OpenAI’s Ada 2 (OpenAI, nd) and the widely-used BERT (Devlin, 2018), transform text into dense vector representations that capture underlying semantic relationships. Dimensionality reduction algorithms like t-SNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), and UMAP (McInnes et al., 2018) are commonly utilized to project the high-dimensional embeddings onto 2D or 3D maps, making complex patterns in the data space more accessible for analysis. While these visualizations primarily focus on interpreting the space of known knowledge, Map2Text extends the capabilities of the map by enabling the generation of new, relevant, and accurately aligned textual content. This shift bridges the gap between data interpretation and creative exploration, offering new possibilities for leveraging data visualizations to facilitate applications such as scientific discovery and innovation.
Controllable Text Generation
Controllable text generation (CTG) refers to the task of generating text according to the given controlled element  (Prabhumoye et al., 2020). It has been a critical focus in recent years and covers a wide range of tasks, such as Attribute-based Generation and Data Augmentation. In Attribute-based Generation, the goal is to generate text that aligns with specific attributes, such as topics, emotions, or keywords, allowing systems to produce interpretable and structured outputs tailored to user-specified needs. For instance, Khalifa et al. (2020) offers a method to manipulate attributes like topic by adjusting latent variables, improving control over the text’s structure and coherence. Similarly, Prabhumoye et al. (2020) examines a broader range of controllable attributes, demonstrating how various methods can enhance the generation process, leading to text that is more aligned with user goals. Data Augmentation is another important task within CTG. Works like Amin-Nejad et al. (2020) and Liu et al. (2020) have highlighted the effectiveness of CTG for data augmentation, where generated text supplements existing data by altering attributes or replacing specific entities, thus broadening the training set and improving model performance. While these works emphasize direct control over attributes such as topic or slot values, Map2Text shifts the focus to leveraging spatial relationship of text representations, using the contextual information in a visualization map to guide content creation.
@@
The evaluation of text generation quality has traditionally relied on similarity-based metrics, broadly categorized into lexical- and semantic-based approaches. These methods have been widely adopted across tasks such as machine translation, summarization, and creative content generation. Lexical-based metrics focus on n-gram overlaps between generated and reference texts. BLEU emphasizes precision, making it a common choice for machine translation (Papineni et al., 2002), while ROUGE, which prioritizes recall, is frequently used for summarization tasks (Lin, 2004). METEOR (Lavie and Agarwal, 2007), another well-known metric, improves by incorporating synonyms and stemming to better handle linguistic variations. However, these metrics often fail to capture the deeper semantic meaning of text. To address these limitations, semantic-based metrics have been developed. Techniques such as BERTScore (Zhang et al., 2019) leverage contextual embeddings from BERT to capture the token-level similarity between sentences. BLEURT (Sellam et al., 2020), by fine-tuning models to predict human judgment, further aligns with human perception and measure beyond surface-level overlaps.
More recently, reasoning-based evaluations have emerged, utilizing the capabilities of LLMs to assess text coherence, consistency, and factual accuracy. Methods like GPTScore (Fu et al., 2023) employ LLM-predicted probabilities as quality scores, while G-Eval (Liu et al., 2023) engages LLMs in tasks such as Likert scale evaluations and pairwise comparisons. These approaches leverage the reasoning abilities of LLMs to evaluate content beyond traditional metrics. FActScore (Min et al., 2023) takes this a step further by breaking generated text into atomic facts and verifying each fact’s support against external knowledge sources. Building on these advancements, our proposed method, Atometric, offers a more comprehensive and interpretable evaluation by assessing both precision and recall of atomic statements at multiple levels of strictness. This approach captures the nuances of semantic relevance and logical coherence of the generated text, addressing gaps left by previous metrics.
@@ 6. Conclusion
In this work, we introduce Map2Text, a novel task for generating new textual content guided by specific positions within a 2D visualization map of a large text corpus. We propose a new evaluation metric, Atometric, to assess the fine-grained alignment of the generated text to gold standard references at varying levels of strictness. Through this metric, we demonstrate the effectiveness of multiple representative candidate methods for Map2Text on datasets in diverse domains. Our findings show the potential of Map2Text as a valuable tool for applications such as scientific idea generation, new persona generation, and LLM red teaming. This initial attempt calls for future research may explore more advanced methods and diverse applications of Map2Text.
6.1. Limitations and Future Work
@@
Novelty of Generated Content:
Our current evaluation does not go beyond measuring similarity between the generated content and a given gold standard reference. For example, it does not consider the novelty of the output or the reference, which is crucial for applications such as research idea generation. One potential solution is to compute the similarity between the generated content and its nearest neighbors (where lower similarity indicates higher novelty) rather than the reference alone. In reality, the quality of the “gold” standard itself is often not guaranteed, as human-generated content can be highly relevant while still lack scientific validity or logical coherence. Future work may employ domain-specific criteria into the evaluation metrics or rely on specific downstream tasks.
Evaluation Limited to Known Positions:
Our evaluation framework is limited to positions on the map for which ground truth content exists but is held out. We do not evaluate the model’s ability to generate content at arbitrary unexplored positions on the map, which would unleash the full potential of Map2Text for open exploration and innovation into the information space. Without the gold stand reference, the quality of the generated new content has to be evaluated otherwise, through online human judgments or the utility of the generated content in downstream tasks, such as the effectiveness of a new red-teaming strategy in testing LLMs.
Optimizing Map2Text Methods:
@@
There is considerable room of improvement beyond the candidate methods tested in this paper. An advanced Map2Text model may leverage global patterns or community structures in the map to guide the generation, beyond the local neighborhood. Another approach is to fine-tune LLMs to utilize spatial tokens that encode positional information directly within their embedding layers. This would allow an LLM to memorize the complex map structure and incorporate spatial context more effectively, while not harming their reasoning ability.
Appendix A Experiment Details
A.1. Results on Extra Metrics
@@
We provide additional experiment results on a broader list of evaluation metrics as a complementary reference for Table 3 and 4.5. The results are shown in Table A.1, 6, and 7.
@@
Persona Atometric Precision (loose) 0.908 0.853 0.822 0.815 0.822 0.758 -
Atometric Precision (moderate) 0.582 0.460 0.449 0.430 0.459 0.365 -
Atometric Precision (strict) 0.201 0.141 0.133 0.125 0.208 0.096 -
Atometric Recall (loose) 0.878 0.850 0.825 0.829 0.822 0.808 -
Atometric Recall (moderate) 0.623 0.533 0.480 0.481 0.467 0.420 -
Atometric Recall (strict) 0.342 0.261 0.229 0.214 0.181 0.148 -
BERTScore Precision 0.893 0.890 0.891 0.891 0.897 0.882 -
BERTScore Recall 0.903 0.900 0.895 0.894 0.890 0.892 -
BLEU (BLEU) 0.075 0.063 0.050 0.060 0.056 0.045 -
Cosine Similarity 0.885 0.879 0.871 0.869 0.868 0.857 -
ROUGE-1 0.360 0.334 0.311 0.309 0.315 0.298 -
ROUGE-L 0.281 0.262 0.244 0.248 0.258 0.231 -
ROUGE-Lsum 0.281 0.262 0.245 0.247 0.258 0.231 -
@@
Table 5. Additional evaluation results of the candidate methods on the Map2Text task. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for strictness level.
Metric P R I C
Atometric Precision (loose) 0.882 0.802 0.453 0.542
Atometric Precision (moderate) 0.548 0.598 0.199 0.103
Atometric Precision (strict) 0.296 0.288 0.103 0.044
Atometric Recall (loose) 0.885 0.764 0.411 0.535
Atometric Recall (moderate) 0.565 0.575 0.184 0.107
Atometric Recall (strict) 0.267 0.279 0.093 0.045
BERTScore Precision 0.898 0.902 0.860 0.864
BERTScore Recall 0.898 0.902 0.855 0.860
BLEU (BLEU) 0.074 0.212 0.020 0.029
Cosine Similarity 0.885 0.925 0.833 0.841
ROUGE-1 0.362 0.475 0.207 0.197
ROUGE-L 0.290 0.361 0.153 0.137
ROUGE-Lsum 0.289 0.361 0.153 0.137
@@
A.2. Human Annotation
Figure 3. Sample screenshot of human annotation protocol.
Appendix B Atometric Details
The evaluation of Atometric involves two steps, decomposition and verification. It is straightforward to adapt Atometric to other datasets with the provided templates. The implementation is available at https://anonymous.4open.science/r/atometric-17A3/. We use gpt-4o-2024-05-13 as the backbone model for all the evaluation.
B.1. Atomic Statement Decomposition
Figure 4. Prompt template for Atometric decomposition. The proposed metric could be adapted to different dataset with minimal change in prompting.
B.2. Atomic Statement Verification
Figure 5. Prompt template for Atometric verification. The proposed metric could be adapted to different dataset with minimal change in prompting.
Appendix C Dataset Details
C.1. Persona
Figure 6. 2D visualization of Persona dataset.
Description
C.2. Red-Teaming Strategies
Figure 7. 2D visualization of Red-Teaming Strategies dataset.
Description
C.3. CS Research Idea & Context
Figure 8. 2D visualization of CS Research Idea dataset.
Figure 9. 2D visualization of CS Research Context dataset.
Description
  • The authors propose a novel approach to geolocalise panoramic images on a 2-D cartographic map by learning a low-dimensional embedded space which allows comparison between an image captured at a location and local neighbourhoods of the map.
  • The authors propose a new approach for predicting SQL query properties, including the query answer size, run-time, and error class, relying on data-driven machine learning techniques and large query workloads, instead of database stats or execution plans.
A few examples of the contexts are shown below:
Figure 10. Prompt for summarization of the Red Teaming LLM dataset.
Appendix D Implementation Details
Repository Access
Our experiment details and code can be found at https://anonymous.4open.science/r/map2text-27FD.
Fine-tuning Experiments Details
We employed the Llama-3.1-70B-Instruct model for fine-tuning, running on 4 NVIDIA A100 GPUs. Each dataset was fine-tuned for three epochs, and for datasets with over 100,000 data points, the process required approximately 50 hours to complete. To enhance the model’s training efficiency, we utilized the Low-Rank Adaptation (LoRA) technique, setting the rank ($r$) to 16 and the scaling factor ($\alpha$) to 32. We set the dropout rate to 0.05 to prevent overfitting, while setting the initial learning rate to 1e-4, weight decay to 0.0, and momentum factor ($\gamma$) to 0.85 to optimize convergence. Besides, mixed precision training was enabled to maximize computational efficiency. For consistency, all experiments were initialized with the same hyperparameter settings. We use temperate $t=1$ during inference.
Embedding Inversion
Prompt Examples
Retrieval Augmented Generation
We use gpt-4o-2024-05-13 for all RAG-based method and set temperature $t=0$ for all generations. Figures 11 provides an example of prompts used in the 1st-order RAG method for the Red Team Attempts dataset in our experiments. For a comprehensive collection of prompts used across all datasets and all methods, please refer to the map2text/model/prompts directory in the repository,which is available at https://anonymous.4open.science/r/map2text-27FD/map2text/model/prompts/.

### v2

@@ 5. Related Work
Visualization of Text Corpora
@@
Low-dimensional data visualization is a powerful tool across various fields, enabling intuitive representation of large datasets (Sadiku et al., 2016). In public health, it aids in tracking disease trends (Kilimba et al., 2015). Renewable energy benefits from visualization for optimizing production and consumption (Kumar and Goyal, 2016), while environmental science relies on geographic visualization for ecological data interpretation (Bohman et al., 2015; MacEachren and Taylor, 2013). Fraud detection employs interactive visualizations to identify anomalies (Dilla and Raschke, 2015), and library management utilizes visualization to optimize resource allocation and strategy (Murphy, 2013). A variety of methods are employed to construct such visualizations. Text embedding models (Muennighoff et al., 2022) such as those proposed in (Li et al., 2023; Wang et al., 2024; Muennighoff et al., 2024) transform the text into dense vector representations that capture underlying semantic relationships. These high-dimensional embeddings are then projected onto lower-dimensional spaces using algorithms like t-SNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), and UMAP (McInnes et al., 2018), enabling the discovery of complex patterns within the data. Alternatively, network-based visualization techniques employ tools such as Graphviz and Gephi (Research, 2008; Bastian et al., 2009) to represent relationships between textual elements. While these visualizations primarily focus on interpreting the space of known knowledge, MapExplorer extends the capabilities of the map by enabling the generation of new, relevant, and accurately aligned textual content. This shift bridges the gap between data interpretation and creative exploration, offering new possibilities for leveraging data visualizations to facilitate applications such as scientific discovery and innovation.
Controllable Text Generation (CTG)
CTG involves generating text under specified constraints or attributes (Prabhumoye et al., 2020). For example, prior work has focused on attribute-based generation (e.g., topic, emotion) by fine-tuning latent variables (Khalifa et al., 2020), and on data augmentation by altering specific entities to enhance training corpora (Amin-Nejad et al., 2020; Liu et al., 2020). While these methods employ explicit signals (e.g., style or slot values) as the “control,” MapExplorer introduces a novel form of controllability by anchoring generation to coordinates in a visualization map. In this setting, the map’s spatial layout dictates the context, allowing users to select points or regions within the embedded space to elicit textual content that reflects local proximity, cluster structure, or overall geometric patterns.
@@
The evaluation of text generation has traditionally relied on similarity-based metrics, broadly categorized into lexical and semantic approaches. Lexical metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Lavie and Agarwal, 2007), measure text quality through n-gram overlaps, capturing surface-level similarity between generated and reference texts. Semantic-based techniques, including BERTScore (Zhang et al., 2019) and BLEURT (Sellam et al., 2020), leverage contextual embeddings or fine-tuned models to better align with human judgments. However, they often lack interpretability and do not provide more informative feedback than a single similarity score. Among reasoning-based metrics, FActScore (Min et al., 2023) stands out by decomposing generated text into atomic facts and verifying their correctness against external knowledge sources. Yet, FActScore exclusively measures factual correctness (precision), without addressing completeness (recall) or accommodating different levels of evaluation strictness given a reference. Motivated by the limitations discussed in Section 3.2, we introduce Atometric, which extends beyond correctness to assess both precision and recall of atomic statements at multiple levels of strictness. By providing a more nuanced and interpretable evaluation, Atometric offers deeper insights into the alignment, coherence, and informativeness of generated text, addressing critical gaps left by prior metrics.
@@ 6. Conclusion
In this work, we introduce MapExplorer, a novel task for generating new textual content guided by specific positions within a 2D visualization map of a large text corpus. We propose a new evaluation metric, Atometric, to assess the fine-grained alignment of the generated text to gold-standard references at varying levels of strictness. Through this metric, we demonstrate the effectiveness of multiple representative candidate methods for MapExplorer on datasets in diverse domains. Our findings show the potential of MapExplorer as a valuable tool for applications such as scientific idea generation, new persona generation, and LLM red teaming. This initial attempt calls for future research to explore more advanced methods and diverse applications of MapExplorer.
6.1. Discussion
@@
Evaluation Scope.
To ensure automatic and scalable evaluation, our framework focuses on measuring the similarity between generated content and a held-out gold-standard reference. While this enables systematic benchmarking, it does not explicitly assess aspects such as novelty or broader utility. Novelty is particularly relevant for applications like research idea generation, but incorporating it into evaluation when a gold standard doesn’t exist remains an open challenge. A promising direction is to measure similarity with nearest neighbors, where lower similarity would indicate higher novelty. Additionally, future work could explore domain-specific evaluation criteria (Si et al., 2024) or assess utility through downstream applications.
The offline evaluation is also limited to querying known positions on the map for which reference content exists rather than testing the model’s ability to generate content at entirely unexplored locations. While this would unlock the full potential of MapExplorer for open-ended discovery, it requires alternative evaluation strategies beyond offline comparison to references, such as online human judgments or downstream task performance.
Moreover, this study primarily aims to introduce and validate the feasibility of MapExplorer rather than to optimize specific methods for this novel task. While we emphasize quantitative and scalable offline evaluation, our tasks and methods are adaptable to online evaluation procedures (Si et al., 2024). We encourage further exploration in these directions, and a live demo is available at https://mapexplorer-app.streamlit.app/ for ad hoc evaluations.
Optimizing MapExplorer Methods:
@@
There is considerable room for improvement beyond the candidate methods tested in this paper. An advanced MapExplorer model may leverage global patterns or community structures in the map to guide the generation beyond the local neighborhood. Another approach is to fine-tune LLMs to utilize spatial tokens that encode positional information directly within their embedding layers. This would allow an LLM to memorize the complex map structure and incorporate spatial context more effectively without harming their reasoning ability. The superior performance of the straightforward RAG + CoT method also implies a promising direction to improve the performance of MapExplorer through optimizing the reasoning chains, potentially through reinforcement learning.
Acknowledgements.
This work is funded in part by the LG AI Research Partnership with the University of Michigan.
Appendix A Interactive Demo
We set up an anonymous demo for MapExplorer, available at
https://mapexplorer-app.streamlit.app/. The demo is developed and deployed using Streamlit and currently supports two medium-sized datasets (Red Teaming Strategies and Persona) along with four prompt-based generation methods due to computational limits.
A sample page of the demo is shown in Figure 3 and 4. The 2D map visualizes the data points using low-dimensional embeddings, with hover functionality to display the corresponding text. To interact with the demo, the users need to first choose the dataset and the generation method. Then. they can either specify a coordinate or choose a random coordinate for generation. The generated content will be displayed at the bottom of the webpage.
Appendix B Experiment Details
B.1. Results on Extra Metrics
@@
We provide additional experiment results on a broader list of evaluation metrics as a complementary reference for Table 4.4 and 4.5. The results are shown in Table E, 6, and 7.
B.2. Human Annotation
Appendix C Atometric Details
The evaluation of Atometric involves two steps, decomposition and verification. It is straightforward to adapt Atometric to other datasets with the provided templates. The implementation is available at https://anonymous.4open.science/r/atometric. We use gpt-4o-2024-05-13 as the backbone model for all the evaluation.
C.1. Atomic Statement Decomposition
C.2. Atomic Statement Verification
Figure 6. Prompt template for Atometric decomposition.
Figure 7. Prompt template for Atometric verification.
Appendix D Dataset Details
D.1. Persona Text
Figure 8. 2D visualization of Persona dataset.
Description
D.2. Red-Teaming Strategies Text
Figure 9. 2D visualization of Red-Teaming Strategies dataset.
Figure 10. Prompt for summarization of the Red Teaming LLM dataset.
Description
D.3. Research Idea & Context Text
Figure 11. 2D visualization of CS Research Idea dataset.
Figure 12. 2D visualization of CS Research Context dataset.
Description
  • The authors propose a novel approach to geolocalise panoramic images on a 2-D cartographic map by learning a low-dimensional embedded space that allows comparison between an image captured at a location and local neighborhoods of the map.
  • The authors propose a new approach for predicting SQL query properties, including the query answer size, run-time, and error class, relying on data-driven machine learning techniques and large query workloads instead of database stats or execution plans.
A few examples of the contexts are shown below:
D.4. Research Context Citation Network
Figure 13. 2D visualization of CS Research Context Citation dataset.
Description
The Research Context Citation Network dataset is constructed by first filtering papers from the OAG v3.1 dataset Zhang et al. (2023) that originate from 17 leading computer science conferences. The selection of these 17 conferences follows the same criteria as in the MASSW paper. Then, to ensure meaningful citation relationships, we remove papers with no citations and those whose citations are isolated, meaning they do not form any connected citation links with other papers within the filtered dataset. For the remaining papers, we follow the same pipeline as MASSW to extract their research context. Using the filtered set of papers and their citation relationships, we construct a directed citation network, where nodes represent papers and directed edges denote citation links. Finally, we apply LargeVis to compute 2D coordinates for each paper, enabling a structured visualization of research contexts based on their citation relationships.
Appendix E Implementation Details
Repository Access
Our experiment details and code can be found at https://anonymous.4open.science/r/mapexplorer.
Fine-tuning Experiments Details
We employed the Llama-3.1-70B-Instruct model for fine-tuning, running on 4 NVIDIA A100 GPUs. Each dataset was fine-tuned for three epochs, and for datasets with over 100,000 data points, the process required approximately 50 hours to complete. To enhance the model’s training efficiency, we utilized the Low-Rank Adaptation (LoRA) technique, setting the rank ($r$) to 16 and the scaling factor ($\alpha$) to 32. We set the dropout rate to 0.05 to prevent overfitting while setting the initial learning rate to 1e-4, weight decay to 0.0, and momentum factor ($\gamma$) to 0.85 to optimize convergence. Besides, mixed precision training was enabled to maximize computational efficiency. For consistency, all experiments were initialized with the same hyperparameter settings. We use temperate $t=1$ during inference.
Embedding Inversion
Retrieval Augmented Generation
We use gpt-4o-2024-05-13 for all RAG-based methods and set temperature $t=0$ for all generations. Figure 14 provides an example of prompts used in the 1st-order RAG method for the Red Team Attempts dataset in our experiments. For a comprehensive collection of prompts used across all datasets and all methods, please refer to the prompts directory in the repository, which is available at https://anonymous.4open.science/r/mapexplorer/map2text/model/prompts/.
@@ Appendix E Implementation Details
Persona (Text) Atometric Precision (loose) 0.908 0.853 0.822 0.815 0.822 0.758 0.914
Atometric Precision (moderate) 0.582 0.460 0.449 0.430 0.459 0.365 0.661
Atometric Precision (strict) 0.201 0.141 0.133 0.125 0.208 0.096 0.378
Atometric Recall (loose) 0.878 0.850 0.825 0.829 0.822 0.808 0.932
Atometric Recall (moderate) 0.623 0.533 0.480 0.481 0.467 0.420 -0.671
Atometric Recall (strict) 0.342 0.261 0.229 0.214 0.181 0.148 0.339
BERTScore Precision 0.893 0.890 0.891 0.891 0.897 0.882 0.901
BERTScore Recall 0.903 0.900 0.895 0.894 0.890 0.892 0.902
BLEU (BLEU) 0.075 0.063 0.050 0.060 0.056 0.045 0.077
Cosine Similarity 0.885 0.879 0.871 0.869 0.868 0.857 0.898
ROUGE-1 0.360 0.334 0.311 0.309 0.315 0.298 0.407
ROUGE-L 0.281 0.262 0.244 0.248 0.258 0.231 0.303
ROUGE-Lsum 0.281 0.262 0.245 0.247 0.258 0.231 0.303
@@
[…]
Atometric Precision (moderate) 0.102 0.072 0.067 0.072 0.047 0.065 0.062
Atometric Precision (strict) 0.033 0.035 0.018 0.017 0.015 0.020 0.026
Atometric Recall (loose) 0.522 0.409 0.430 0.444 0.294 0.424 0.503
Atometric Recall (moderate) 0.070 0.048 0.047 0.034 0.031 0.077 0.077
Atometric Recall (strict) 0.019 0.016 0.016 0.005 0.009 0.023 0.018
BERTScore Precision 0.866 0.867 0.869 0.869 0.858 0.807 0.860
BERTScore Recall 0.858 0.854 0.854 0.852 0.849 0.820 0.856
BLEU (BLEU) 0.014 0.010 0.009 0.007 0.010 0.007 0.011
Cosine Similarity 0.827 0.818 0.816 0.814 0.803 0.811 0.831
ROUGE-1 0.175 0.159 0.164 0.153 0.155 0.161 0.192
ROUGE-L 0.126 0.113 0.115 0.111 0.106 0.100 0.126
ROUGE-Lsum 0.127 0.113 0.115 0.112 0.106 0.116 0.126
Table 5. Additional evaluation results of the candidate methods on the MapExplorer task. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for strictness level.
Metric Persona (Text) Red Teaming (Text) Idea (Text) Context (Text) Context Citation (Network)
Atometric Precision (loose) 0.882 0.802 0.453 0.542 0.459
Atometric Precision (moderate) 0.548 0.598 0.199 0.103 0.095
Atometric Precision (strict) 0.296 0.288 0.103 0.044 0.032
Atometric Recall (loose) 0.885 0.764 0.411 0.535 0.431
Atometric Recall (moderate) 0.565 0.575 0.184 0.107 0.066
Atometric Recall (strict) 0.267 0.279 0.093 0.045 0.021
BERTScore Precision 0.898 0.902 0.860 0.864 0.856
BERTScore Recall 0.898 0.902 0.855 0.860 0.854
BLEU (BLEU) 0.074 0.212 0.020 0.029 0.015
Cosine Similarity 0.885 0.925 0.833 0.841 0.820
ROUGE-1 0.362 0.475 0.207 0.197 0.166
ROUGE-L 0.290 0.361 0.153 0.137 0.113
ROUGE-Lsum 0.289 0.361 0.153 0.137 0.114